OpenStudy (exoexoexo):

find the x and y intercept of y=(3/5)x+5

OpenStudy (texaschic101):

In y = mx + b form, the y intercept is in the b position. y = mx + b y = 3/5x + 5.....so the y intercept is 5 or (0,5) to find the x intercept, sub in 0 for y y = 3/5x + 5 0 = 3/5x + 5 -3/5x = 5 x = 5/(-3/5) x = 5 * - 5/3 x = - 25/3....so the x intercept is (-25/3,0)